OpenStudy (anonymous):

Wynonna opened a savings account 30 years ago with a deposit of $5,140.17. The account has an interest rate of 3.6% compounded quarterly. How much interest has Wynonna earned?

OpenStudy (anonymous):

A = P(1+ (r/n))^nt Where P is the Principle, r is the rate, n is how often it's compounded, and t is the time in years. A = 5,140.17 * (1 + (.036/4))^4*30 5,140.17 * (1.009)^120 A = 15,063.21

OpenStudy (anonymous):

15,063.21 is your answer

OpenStudy (anonymous):

I got A=5,140.16(1+.036/4)^120 and got 16,598.84 as my ending answer which isnt one of the options

OpenStudy (anonymous):

ohhh okay i see what i did wrong, thank you dakota!

OpenStudy (anonymous):

15,063.21 is the correct answer. Can I have a medal plz?

OpenStudy (anonymous):

thnx